2. Thatcher Corporation’s bonds will mature in 20 years. The bonds have a face value of $1,000 and an 9% coupon rate, paid semiannually. The price of the bonds is $1,350. The bonds are callable in 5 years at a call price of $1,250. What is their yield to maturity? What is their yield to call?
Answer #1

1)

Number of periods = 20 * 2 = 40

Coupon = (0.09 * 1000) / 2 = 45

Yield to maturity = 5.98%

Keys to use in a financial calculator:

2nd I/Y 2

FV 1000

PV -1,350

PMT 45

N 40

CPT I/Y

2)

Number of periods = 5 * 2 = 10

Coupon = (0.09 * 1000) / 2 = 45

Yield to call = 5.36%

Keys to use in a financial calculator:

2nd I/Y 2

FV 1,250

PV -1,350

PMT 45

N 10

CPT I/Y